OpenStudy (rezper):

What is a? 5/a - 2/ a - 1 = 2/ a - 1

OpenStudy (3mar):

Your equation is \[\frac{ 5 }{ a } -\frac{ 2 }{ a } = \frac{ 2 }{ a }-1\]

OpenStudy (rezper):

How did you right it like that?

OpenStudy (3mar):

You don't know?

OpenStudy (rezper):

Also that isnt it

OpenStudy (rezper):

Nope but Ill write it out if you tell me :D

OpenStudy (3mar):

It is easy, man!

OpenStudy (rezper):

How?

OpenStudy (rezper):